A student opens a math book to two facing pages. The product of the page is 156. Find the page numbers.
The first page is ____.
The second page is ____.

Answer by jorel1380(3719)

n(n+1)=156
n2+n-156=0
(n+13)(n-12)=0
n=12 or -13
The two adjoining pages are pages # 12 and 13..